1. A share of common stock will pay $1.00 in a year, and at that time, the stock is sold for $50. If the rate of return (Rs) is 10%, then what is the price of the stock?

Stock Price Today, P0 = ________

2. A share of common stock will pay $1.00 in a year, $3.00 in two years, and at that time, the stock is sold for $50. If the rate of return (Rs) is 10%, then what is the price of the stock?

Stock Price Today, P0 = ________

3. A share of common stock just paid $1.00, for the next three years the dividend will grow by 10%, and in the fourth year, the dividend will start to grow continuously at 5% forever. If the rate of return (Rs) is 10%, then what is the price of the stock?

Stock Price Today, P0 = ________
Answer #1

1.Price of stock=Future dividend and value*Present value of discounting factor(rate%,time period)

=1/1.1+50/1.1

=$46.36(Approx).

2.Price of stock=Future dividend and value*Present value of discounting factor(rate%,time period)

=1/1.1+3/1.1^2+50/1.1^2

=$44.71(Approx).

3.

D1=(1*1.1)=1.1

D2=(1.1*1.1)=1.21

D3=(1.21*1.1)=1.331

Value after year 3=(D3*Growth Rate)/(Required rate-Growth Rate)

=(1.331*1.05)/(0.1-0.05)

=$27.951

Hence price of stock=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.1/1.1+1.21/1.1^2+1.331/1.1^3+$27.951/1.1^3

=$24